OpenStudy (anonymous):

the gradient function of a curve is px^2-qx,where p and q are constant .the curve has a turning point at (2,-1).the gradient of the tangent to the curve at the point x=-1 is 9.find the value of p and q.

OpenStudy (callisto):

Can you differentiate y = px^2-qx with respect to x first?

OpenStudy (anonymous):

yes.. but why ? ><

OpenStudy (callisto):

Because the words "turning point" imply that the first derivative is 0 at (2, -1) And "gradient of tangent at x=-1" is the first derivative of the function at x=-1. So, by finding the first derivative of px^2-qx, we can set two equations, so that we can solve p and q.

OpenStudy (anonymous):

can you show me the step ?

OpenStudy (callisto):

Do you know how to differentiate px^2-qx?

OpenStudy (anonymous):

yes. dy/dx= 2p-q. isit ?

OpenStudy (callisto):

It... is not. Check the first term again.

OpenStudy (anonymous):

2px-q.

OpenStudy (callisto):

Yup! So, from " a turning point at (2,-1)" Just sub x=2 and y=-1 into y=px^2-qx. This is the first equation you get. What would that be?

OpenStudy (anonymous):

4p-2q=-1 -----(1)

OpenStudy (callisto):

Yes. Next, for " tangent to the curve at the point x=-1 is 9", we put dy/dx=9 and x=-1 into the first derivative. That is our second equation. What is that?

OpenStudy (callisto):

* for "the gradient of the tangent to the curve at the point x=-1 is 9"

OpenStudy (anonymous):

9=-2p-q . ---(2) thn use similtaneous solve it right ?

OpenStudy (callisto):

Exactly :)

OpenStudy (anonymous):

thanks alot ! :)

OpenStudy (callisto):

Welcome~